How do you solve this question. 1/8-3/4x=1/16
Answer by benazir.sj@gmail.com(109)

1/8-3/4x=1/16
solution:
1/8 -3/4x = 1/16
-3/4x =1/16-1/8
-3/4x =1/16-1*2/8*2
-3/4x=1-2/16
-3/4x=-1/16
-3/4x *-4/3 = -1/16*(-4/3)
x=4/(16*3)
x=1/4*3
x=1/12